# Quantum Steenrod operations of symplectic resolutions

### v2

@@ 2.5.1. Quantum Steenrod operations for the $\infty$-fiber
with incidence constraints under evaluation maps $\mathcal{M}^{eq}_{\beta}\to E({\mathbb{Z}/p}\times{T})\times(X\times X^{p}\times X)$. Consider the action (cf. (2.10))
@@
where $\rho_{\beta}$ acts diagonally. The action $\sigma_{X}^{\beta}$ is defined so that $(\mathrm{id},\phi_{\beta}):(X\times X^{p}\times X;\rho\circ\sigma_{X})\to(X\times X^{p}\times X;\sigma_{X}^{\beta})$ is a $({\mathbb{Z}/p}\times T)$-equivariant map. Under $\sigma_{X}^{\beta}$, the evaluation map becomes equivariant and one can define the moduli spaces with incidence constraints to define the structure constants of $\Sigma_{b}^{\beta}$. The following observation shows that the twisted quantum Steenrod operations are nothing but conjugation of the ordinary quantum Steenrod operations by the twisted-linear isomorphism $\Phi_{\beta}$.
@@
{tikzcd}
@@ 2.5.2. A TQFT argument
It is convenient to introduce the following family of curves. Fix $S=\mathbb{P}^{1}$, and take the trivial family of curves $C\times S\to S$. Let $\mathcal{C}$ be the blowup of $C\times S$ at two points $(0,0)$, (,)($\infty$,$\infty$)( ∞ , ∞ ). The resulting family $\mathcal{C}\to S$ of nodal genus 00 curves over $S$ has two nodal fibers at $0,\infty\in S$ and smooth fibers at all other values of $v\in S$. Let $z_{*}:S\to\mathcal{C}$ be a section of this family defined as the proper transform of the diagonal section. Also fix $z_{0},z_{\infty}:S\to\mathcal{C}$ the sections defined as the proper transform of the constant sections $\{0\}\times S$, {}×$\{\infty\}\times S$. Let $L\subseteq S$ be the positive real line compactified by $0,\infty\in S$, oriented from 00 to $\infty$.
Fix the identification $\mathcal{C}_{v}\cong C$ for $v\neq 0,\infty$. For $v=0$ and $v=\infty$, the fiber $\mathcal{C}_{v}$ is a nodal genus 00 curves with two components $\mathcal{C}_{v,+}\cong C$ and $\mathcal{C}_{v,-}$. We moreover fix the biholomorphism of the exceptional curve $\mathcal{C}_{0,-}\cong C$, so that $0,1,\infty\in C$ corresponds to $z_{0}(0)$, $z_{*}(0)$, and the nodal point respectively. Similarly, fix the biholomorphism $\mathcal{C}_{\infty,-}\cong C$, so that $0,1,\infty\in C$ corresponds to the nodal point, $z_{*}(\infty)$, $z_{\infty}(\infty)$ respectively. For $z\in\mathcal{C}$, denote by $\hat{z}$ the image under projection $\mathcal{C}\to C$; if $z\in\mathcal{C}_{v,-}$ for $v=0$ or $v=\infty$, then $\hat{z}=v$, but otherwise $\hat{z}={z}$ under the identification $C\cong\mathcal{C}_{v}$. Denote the complement in $\mathcal{C}$ of the two nodal points by $\mathcal{C}^{\mathrm{reg}}$.
@@
(2.70) $\mathcal{N}^{eq}_{L}(E_{\beta})=\left\{v\in L,\ u:\mathcal{C}_{v}\to E_{\beta},\ (w,w^{\prime})\in E({\mathbb{Z}/p}\times{T}):\pi_{E}(u(z))=\hat{z},\ \overline{\partial}_{\widetilde{J}}u=\nu_{\mathcal{C}_{v},w,w^{\prime}}^{eq}\right\}$
@@
(2.74) $\Phi_{\beta}\circ S_{\beta}\circ\Sigma_{b}=\Sigma_{b}\circ\Phi_{\beta}\circ S_{\beta}.$
@@ 3.1.1. Geometry and topology
Let $G$ be a complex semisimple connected Lie group and $B$ be a Borel subgroup, and $T_{\mathbb{C}}$ be the maximal torus in $B$. Correspondingly consider the Lie algebras $\mathfrak{g}$, $\mathfrak{b}$, $\mathfrak{h}$. Using the Killing form, we will identify $\mathfrak{g}^{*}\cong\mathfrak{g}$ interchangeably. Note that $G$ acts on the set of Borel subalgebras of $\mathfrak{g}$, with $B$ being the stabilizer of $\mathfrak{b}$. Therefore the set of Borel subalgebras is a homogeneous space $G/B$, the flag variety. The flag variety is a smooth Fano variety, hence a monotone symplectic manifold by the anticanonical polarization. The group $G$, and the maximal torus subgroup $T_{\mathbb{C}}$, naturally acts on $G/B$.
The holomorphic cotangent bundle $X=T^{*}(G/B)$ of the flag variety is a holomorphic symplectic manifold equipped with the holomorphic Liouville form $\omega_{\mathbb{C}}$. By restriction from the projectivization $\mathbb{P}(T^{*}G/B\oplus\underline{\mathbb{C}})\to G/B$, the cotangent bundle $T^{*}(G/B)$ admits a Kähler structure $\omega$ compatible with its integrable complex structure. The Kähler form $\omega$ is indeed non-exact, with the zero section $G/B$ being a symplectic submanifold, and $T^{*}(G/B)$ with its integrable complex structure does admit many holomorphic curves. This is in contrast with the symplectic structure such as $\mathrm{Re}(\omega_{\mathbb{C}})$, for which the zero section $G/B$ is a Lagrangian submanifold.
@@
There is also a complete hyperKähler metric on $X$ such that the metric restricts to the homogeneous metric on the zero section $G/B$ and is compatible with the holomorphic Liouville form $\omega_{\mathbb{C}}$ [Nak94], [Biq96]. ∎
@@
(3.1) $T^{*}(G/B)\cong G\times^{B}\mathfrak{n}=\{(gB,x):g^{-1}xg\in\mathfrak{n}\}\subseteq G/B\times\mathcal{N}$
@@
(3.2) $H^{*}_{\mathbf{T}}(\mathrm{pt})=\mathbb{F}_{p}[\![\lambda,h]\!]$
where $\lambda=\lambda_{1},\dots,\lambda_{r}$ and $h$ denote the equivariant parameters for the $T$ and $S^{1}$-action, respectively.
@@
It is essential to work equivariantly with respect to the $S^{1}$-action rotating the fibers for Gromov–Witten theory to be nontrivial for $X$. The ordinary Gromov–Witten theory for $X$ is trivial because $X$ admits a deformation to an affine variety, by Grothendieck’s simultaneous resolution. Such deformation cannot be chosen in a way preserving the $S^{1}$-symmetry, and accordingly the (reduced) Gromov–Witten theory may be nontrivial, see [BMO11]. ∎
@@
The flag variety $G/B$ admits a complex cell decomposition [CG97, Corollary 3.1.12] into Schubert cells $\mathring{X}_{w}=BwB/B$ indexed by elements of the Weyl group $w\in W=N_{G}(T)/T$. The cells $\mathring{X}_{w}$ are affine spaces of dimension $\ell(w)$ for the length function $\ell:W\to\mathbb{Z}_{\geq 0}$. The closures $X_{w}=\overline{\mathring{X}_{w}}$ are the Schubert varieties which are themselves union of Schubert cells, $X_{w}=\bigsqcup_{v\leq w}\mathring{X}_{v}$ for $v\leq w$ in a partial order on $W$ known as the Bruhat order. The fundamental classes $[X_{w}]$ yield an additive basis of cohomology $H^{*}(G/B)$. In particular, the cohomology $H^{*}(G/B;\mathbb{Z})$ is concentrated in even degrees, is torsion-free, and the classes are represented by genuine cycles (cf. 2.12).
@@ 3.1.2. Cohomological stable envelopes
The cycles $Y_{w}$ are holomorphic Lagrangian cycles of $X=T^{*}(G/B)$, which are manifestly $\mathbf{T}$-invariant hence representing equivariant cohomology classes in $H^{*}_{\mathbf{T}}(X)$. Denote by $\mathbb{Y}:=\bigcup_{w\in W}Y_{w}$ the union of all conormals.
Definition 3.5.
The conormals at infinity $\mathfrak{f}_{+}$ is the union of ideal boundary of conormals, that is for $\mathbb{Y}=\bigcup_{w\in W}Y_{w}$,
(3.3) $\mathfrak{f}_{+}:=\partial_{\infty}\mathbb{Y}=\bigcup_{w\in W}N^{*}_{\infty}\mathring{X}_{w}.$
We define the ideal boundary as the boundary of the intersection of $\mathbb{Y}$ with a sufficiently small radially symmetric neighborhood of the section at infinity of the projective compactification $\mathbb{P}(T^{*}G/B\oplus\underline{\mathbb{C}})\to G/B$. We may consider $\mathfrak{f}_{+}$ as a closed subset of $X$ by choosing such a sufficiently small neighborhood. The homeomorphism type of the ideal boundary is well-defined regardless of the choice by the $\mathbb{C}^{*}$-action dilating the cotangent fibers.
@@
The conormal cycles $[Y_{w}]$ span the Poincaré dual of middle-dimensional relative cohomology $H^{\dim_{\mathbb{C}}(X)}(X,\mathfrak{f}_{+})^{\vee}$ of rank $|W|$.
@@
The conormal cycles $[Y_{w}]$ represent classes in the Borel–Moore homology of $\mathbb{Y}:=\bigcup_{w\in W}Y_{w}$, hence classes in the dual $H_{*}(\mathbb{Y})^{\vee}$ using the pairing between $H_{*}(\mathbb{Y})$ and $H_{*}^{\mathrm{BM}}(\mathbb{Y})$. By Poincaré duality, $H_{*}(\mathbb{Y})\cong H^{*}_{c}(\mathbb{Y})\cong H^{*}(\mathbb{Y},\mathfrak{f}_{+})\cong H^{*}(X,\mathfrak{f}_{+})$ where the last isomorphism is induced by the homotopy equivalence of pairs.
Choose a filtration of $\mathbb{Y}$ into locally closed subsets $\emptyset=Y_{0}\subseteq Y_{1}\subseteq\cdots\subseteq Y_{|W|}=\mathbb{Y}$ such that every $Y_{j}\setminus Y_{j-1}$ is a conormal of a Schubert cell $N^{*}\mathring{X}_{w}$. Note that the compactly supported cohomology $H^{*}_{c}(Y_{j}\setminus Y_{j-1})$ is of rank $1$ and concentrated in degree $\dim_{\mathbb{C}}(X)=\frac{1}{2}\dim_{\mathbb{R}}(X)$, spanned by the Poincaré dual of $[N^{*}\mathring{X}_{w}]$. Iterative application of the long exact sequence for compactly supported cohomology implies that $H^{*}(X,\mathfrak{f}_{+})$ is concentrated in degree $\dim_{\mathbb{C}}(X)$ and is of rank $|W|$. Since for each $w\in W$, $Y_{w}=N^{*}\mathring{X}_{w}\cup\bigsqcup_{v<w}N^{*}\mathring{X}_{v}$, it follows that the Poincaré duals of $[Y_{w}]$ are linearly independent in $H^{*}(X,\mathfrak{f}_{+})$, as desired. ∎
Note that one can also start from the anti-dominant cocharacter and correspondingly define the opposite Schubert cells $\mathring{X}^{w}$, the opposite conormals $Y^{w}$, and the opposite conormals at infinity $\mathfrak{f}_{-}$. 3.4 implies that the intersection pairing between $H^{*}(X,\mathfrak{f}_{+})$ and $H^{*}(X,\mathfrak{f}_{-})$ is upper-triangular in the basis of conormals $Y_{w}$ and $Y^{w}$, with diagonal entries given by ±$\pm 1$. The equivariant cohomology classes in $H^{*}_{\mathbf{T}}(X)$ represented by the cycles $Y_{w}$ therefore factor through the map
(3.4) $H^{\dim_{\mathbb{C}}(X)}(X,\mathfrak{f}_{+})^{\vee}\cong H^{\dim_{\mathbb{C}}(X)}(X,\mathfrak{f}_{-})\to H^{*}_{\mathbf{T}}(X,X\setminus\mathbb{Y})\to H^{*}_{\mathbf{T}}(X)$
where the second map uses the retract $(X,X\setminus\mathbb{Y})\to(X,\mathfrak{f}_{-})$ induced by the dominant $\mathbb{C}^{*}$-action (induced by complexifying the choice of dominant cocharacter $\beta:S^{1}\to T$) and the last map is a part of the exact sequence of pairs.
@@
To define the stable bases, note that the cotangent fiber at any $T$-fixed point $wB/B\in G/B$ is a $T$-representation with Euler class $e_{w}=e^{T}(T^{*}_{wB/B}G/B)$. For $N_{w}=T_{(wB/B,0)}(X)$ the tangent space of $(wB/B,0)$ in $X=T^{*}(G/B)$, consider the ${\mathbf{T}}$-invariant decomposition $N_{w}=N_{w,+}\oplus N_{w,-}$ into positive and negative $T$-weights (for our choice of dominant $\beta:S^{1}\to T$). By the duality from holomorphic symplectic form, one concludes that $N_{w,+}\cong N_{w,-}^{\vee}$ up to a twist of a character of the $S^{1}$-action rotating the holomorphic symplectic form. Fix the sign ±$\pm e(N_{w,-})$ so that
@@
satisfying the following properties. Denoting $w\in W$ to represent the fundamental class in $H^{0}_{\mathbf{T}}(X^{T})$ of the isolated fixed point $(wB/B,0)\in X^{T}$, (i) $\mathrm{Stab}_{+}(w)|_{H^{*}_{{\mathbf{T}}}(X\setminus\mathbb{Y})}=0$, (ii) $\mathrm{Stab}_{+}(w)|_{w}=\pm e(N_{w,-})$ where the sign is determined by (3.5), (iii) $\mathrm{Stab}_{+}(w)|_{v}$ is divisible by $h$ for $v<w$ in the Bruhat order.
@@
Property (i) with 3.6 implies that the stable envelope classes $\mathrm{Stab}_{+}(w)$ for $w\in W$ are linear combinations of the conormal cycles $Y_{w}$ in $H^{*}_{\mathbf{T}}(X)$. Property (ii) implies that the stable envelope classes form a basis in $H^{*}_{\mathbf{T}}(X)_{\mathrm{loc}}$ after localization.
Moreover, the stable envelopes are defined via Lagrangian correspondences, and hence takes middle degree to middle degree. As $w\in W$ has a corresponding class $[(wB/B,0)]\in H^{0}_{\mathbf{T}}(X^{T})$, the stable envelope class $\mathrm{Stab}_{+}(w)\in H^{\dim_{\mathbb{C}}X}_{\mathbf{T}}(X)$ lies in middle dimension (which is also evident from property (ii)). ∎

### v3

@@ 2.5.1. Quantum Steenrod operations for the $\infty$-fiber
with incidence constraints under evaluation maps $\mathcal{M}^{eq}_{\beta}\to E({\mathbb{Z}/p}\times T)\times(X\times X^{p}\times X)$. Consider the action (cf. (2.10) where $\sigma_{X}$ is defined as the cyclic permutation action on $X^{p}$)
@@
where $\rho_{\beta}$ acts diagonally. The action $\sigma_{X}^{\beta}$ is defined so that $(\mathrm{id},\phi_{\beta}):(X\times X^{p}\times X;\rho\circ\sigma_{X})\to(X\times X^{p}\times X;\sigma_{X}^{\beta})$ is a $({\mathbb{Z}/p}\times T)$-equivariant map. Using the action $\sigma_{X}^{\beta}$, the evaluation map becomes equivariant and one can define the moduli spaces with incidence constraints to define the structure constants of $\Sigma_{b}^{\beta}$. The following observation shows that the twisted quantum Steenrod operations are nothing but conjugation of the ordinary quantum Steenrod operations by the twisted-linear isomorphism $\Phi_{\beta}$.
@@
${H^{*}_{{\mathbb{Z}/p}\times T}(X;\Lambda|\rho_{\beta})}$${H^{*}_{{\mathbb{Z}/p}\times T}(X;\Lambda|\rho_{\beta})}$${H^{*}_{{\mathbb{Z}/p}\times T}(X;\Lambda|\rho_{0})}$${H^{*}_{{\mathbb{Z}/p}\times T}(X;\Lambda|\rho_{0})}$$\scriptstyle{\Phi_{\beta}}$$\scriptstyle{\Sigma_{b}^{\beta}}$$\scriptstyle{\Phi_{\beta}}$$\scriptstyle{\Sigma_{b}}$
@@ 2.5.2. A TQFT argument
It is convenient to introduce the following family of curves. Fix $S=\mathbb{P}^{1}$, and take the trivial family of curves $C\times S\to S$ where $C$ is our distinguished parameterized copy of $\mathbb{P}^{1}$ which is the domain of the maps we are counting. Let $\mathcal{C}$ be the blowup of $C\times S$ at two points $(0,0)$, (,)($\infty$,$\infty$)( ∞ , ∞ ). The resulting family $\mathcal{C}\to S$ of nodal genus 00 curves over $S$ has two nodal fibers at $0,\infty\in S$ and smooth fibers at all other values of $v\in S$. Let $z_{*}:S\to\mathcal{C}$ be a section of this family defined as the proper transform of the diagonal section. Also fix $z_{0},z_{\infty}:S\to\mathcal{C}$ the sections defined as the proper transform of the constant sections $\{0\}\times S$, {}×$\{\infty\}\times S$. Let $L\subseteq S$ be the positive real line compactified by $0,\infty\in S$, oriented from 00 to $\infty$.
Fix the identification $\mathcal{C}_{v}\cong C$ for $v\neq 0,\infty$. For $v=0$ and $v=\infty$, the fiber $\mathcal{C}_{v}$ is a nodal genus 00 curves with two components $\mathcal{C}_{v,+}\cong C$ and $\mathcal{C}_{v,-}$. We moreover fix the biholomorphism of the exceptional curve $\mathcal{C}_{0,-}\cong C$, so that $0,1,\infty\in C$ corresponds to $z_{0}(0)$, $z_{*}(0)$, and the nodal point respectively. Similarly, fix the biholomorphism $\mathcal{C}_{\infty,-}\cong C$, so that $0,1,\infty\in C$ corresponds to the nodal point, $z_{*}(\infty)$, $z_{\infty}(\infty)$ respectively. For $z\in\mathcal{C}$, denote by $\hat{z}$ the image under projection $\mathcal{C}\to C\times S\to C$; if $z\in\mathcal{C}_{v,-}$ for $v=0$ or $v=\infty$, then $\hat{z}=v$, but otherwise $\hat{z}={z}$ under the identification $C\cong\mathcal{C}_{v}$. Denote the complement in $\mathcal{C}$ of the two nodal points by $\mathcal{C}^{\mathrm{reg}}$.
@@
(2.72) $\mathcal{N}^{eq}_{L}(E_{\beta})=\left\{v\in L,\ u:\mathcal{C}_{v}\to E_{\beta},\ (w,w^{\prime})\in E({\mathbb{Z}/p}\times T):\pi_{\beta}(u(z))=\hat{z},\ \overline{\partial}_{\widetilde{J}}u=\nu_{\mathcal{C}_{v},w,w^{\prime}}^{eq}\right\}$
@@
(2.76) $\Sigma_{b}\circ\Phi_{\beta}\circ S_{\beta}=\Phi_{\beta}\circ S_{\beta}\circ\Sigma_{b}.$
@@ 3.1.1. Geometry and topology
Let $G$ be a complex semisimple connected Lie group and $B$ be a Borel subgroup, and $T_{\mathbb{C}}$ be the maximal torus in $B$. Correspondingly consider the Lie algebras $\mathfrak{g}$, $\mathfrak{b}$, $\mathfrak{h}$. Using the Killing form, we will identify $\mathfrak{g}^{*}\cong\mathfrak{g}$ interchangeably. Note that $G$ acts transitively on the set of Borel subalgebras of $\mathfrak{g}$ by conjugation, with $B$ being the stabilizer of $\mathfrak{b}$. Therefore the set of Borel subalgebras is a homogeneous space $G/B$, the flag variety. The flag variety is a smooth Fano variety, hence a monotone symplectic manifold by the anticanonical polarization. The group $G$, and the maximal torus subgroup $T_{\mathbb{C}}$, naturally acts on $G/B$.
The holomorphic cotangent bundle $X=T^{*}(G/B)$ of the flag variety is a holomorphic symplectic manifold equipped with the standard holomorphic symplectic form $\omega_{\mathbb{C}}$. By restriction from the projectivization $\mathbb{P}(T^{*}G/B\oplus\underline{\mathbb{C}})\to G/B$, the cotangent bundle $T^{*}(G/B)$ also admits a Kähler structure $\omega$ compatible with its integrable complex structure. The Kähler form $\omega$ is indeed non-exact, with the zero section $G/B$ being a symplectic submanifold, and $T^{*}(G/B)$ with its integrable complex structure does admit many holomorphic curves. This is in contrast with the symplectic structure such as $\mathrm{Re}(\omega_{\mathbb{C}})$, for which the zero section $G/B$ is a Lagrangian submanifold.
@@
There is also a complete hyperKähler metric on $X$ such that the metric restricts to the homogeneous metric on the zero section $G/B$ and is compatible with the standard holomorphic symplectic form $\omega_{\mathbb{C}}$ [Nak94], [Biq96]. ∎
@@
(3.1) $T^{*}(G/B)\cong G\times^{B}\mathfrak{n}:=\{(gB,x):g^{-1}xg\in\mathfrak{n}\}\subseteq G/B\times\mathcal{N}$
@@
(3.2) $H^{*}_{\mathbf{T}}(\mathrm{pt})=\mathbb{F}_{p}[\![\lambda,h]\!]:=\mathbb{F}_{p}[\![\lambda_{1},\dots,\lambda_{r},h]\!]$
where $\lambda$ is a shorthand for the collection of all equivariant parameters $\lambda_{1},\dots,\lambda_{r}$ for the $T$-action, and $h$ denotes the equivariant parameter for the rotating $S^{1}$-action.
@@
It is essential to work equivariantly with respect to the $S^{1}$-action rotating the fibers for Gromov–Witten theory to be nontrivial for $X$. The ordinary Gromov–Witten theory for $X$ is trivial because $X$ admits a deformation to an affine variety, by Grothendieck’s simultaneous resolution. Such deformation cannot be chosen in a way preserving the $S^{1}$-symmetry, and accordingly the (reduced) Gromov–Witten theory may be nontrivial, see [BMO11, Section 4.2]. ∎
Example 3.4.
For $G=\mathrm{SL}_{2}(\mathbb{C})$, we may take the subgroup $B$ of upper triangular matrices, with $T_{\mathbb{C}}$ the diagonal matrices. The homogeneous space $G/B$ is isomorphic to $\mathbb{P}^{1}$. On the level of Lie algebras, $\mathfrak{g}=\mathfrak{sl}_{2}=\{\begin{pmatrix}a&b\\ c&-a\end{pmatrix}\}$ is the Lie algebra of traceless matrices, and the nilpotent cone $\mathcal{N}$ can be identified with matrices with vanishing determinant $a^{2}+bc=0$.
In particular, the nilpotent cone $\mathcal{N}\cong\{a^{2}+bc=0\}$ is a copy of a singular quadric cone in ambient dimension 3. Its resolution of singularity is given by the blowup of the singular point, which is identified with $T^{*}\mathbb{P}^{1}$; the zero section $\mathbb{P}^{1}$ corresponds to the exceptional divisor of the blowup.
There is a $\mathbf{T}=S^{1}\times S^{1}$ action on the resolution $T^{*}\mathbb{P}^{1}$; the first factor $T=S^{1}$ acts on the base by rotating the $\mathbb{P}^{1}$ along a fixed axis extended to an action on the cotangent bundle, and the latter factor $S^{1}$ acts on $T^{*}\mathbb{P}^{1}$ by rotating the cotangent fibers.
@@
The flag variety $G/B$ admits a complex cell decomposition [CG97, Corollary 3.1.12] into Schubert cells $\mathring{X}_{w}=BwB/B$ indexed by elements of the Weyl group $w\in W=N_{G}(T)/T$. The cells $\mathring{X}_{w}$ are affine spaces of complex dimension $\ell(w)$ for the length function $\ell:W\to\mathbb{Z}_{\geq 0}$. The closures $X_{w}=\overline{\mathring{X}_{w}}$ are the Schubert varieties which are themselves union of Schubert cells, $X_{w}=\bigsqcup_{v\leq w}\mathring{X}_{v}$ for $v\leq w$ in a partial order on $W$ known as the Bruhat order. The fundamental classes $[X_{w}]$ yield an additive basis of cohomology $H^{*}(G/B)$. In particular, the cohomology $H^{*}(G/B;\mathbb{Z})$ is concentrated in even degrees, is torsion-free, and the classes are represented by genuine cycles (cf. 2.2).
@@
Example 3.6.
In the case of $T^{*}\mathbb{P}^{1}$, the Weyl group $W$ is identified with $\mathbb{Z}/2$ with $1\in W$ and the nontrivial element $s\in W$. The Bruhat order is given by $1<s$. The Weyl group elements are represented in $G=\mathrm{SL}_{2}(\mathbb{C})$ by permutation matrices.
Correspondingly there are two fixed points of the $T=S^{1}$-action on $G/B$, given by the cosets $B/B$ and $sB/B$. Under the isomorphism $G/B\cong\mathbb{P}^{1}$, these correspond to the two fixed points of a rotation $S^{1}$-action (the north and south poles). Fixing a dominant cocharacter $\beta:S^{1}\to T=S^{1}$ (in this case, this just means that the degree of $\beta$ is positive, the corresponding Schubert cells are given by $\mathring{X}_{1}$ = the north pole (a 00-cell), and $\mathring{X}_{s}$ = the complement of the north pole (a $2$-cell). The corresponding Schubert varieties are $X_{1}=\mathring{X}_{1}$ = the north pole, and $X_{s}=\mathring{X}_{s}\cup\mathring{X}_{1}$ = the whole $\mathbb{P}^{1}$.
Opposite Schubert stratification is given by reversing the cocharacter $\beta$, so we have $\mathring{X}^{1}$ = the complement of the south pole, $\mathring{X}^{s}$ = the south pole. Indeed, since the two poles are disjoint, $X_{1}\cap X^{s}=\emptyset$.
@@ 3.1.2. Cohomological stable envelopes
The (locally finite) cycles $Y_{w}$ are holomorphic Lagrangian (locally finite) cycles of $X=T^{*}(G/B)$, which are manifestly $\mathbf{T}$-invariant hence representing equivariant cohomology classes in $H^{*}_{\mathbf{T}}(X)$. Denote by $\mathbb{Y}:=\bigcup_{w\in W}Y_{w}$ the union of all conormals.
Note that one can also start from the anti-dominant cocharacter and correspondingly define the opposite Schubert cells $\mathring{X}^{w}$, the opposite conormals $Y^{w}$, and the union of opposite conormals $\mathbb{Y}^{-}=\bigcup_{w\in W}Y^{w}$. 3.5 implies that the intersection pairing between $H^{*}(X,X\setminus\mathbb{Y})$ and $H^{*}(X,X\setminus\mathbb{Y}^{-})$ is upper-triangular in the basis of (Poincaré duals of) conormals $Y_{w}$ and $Y^{w}$, with diagonal entries given by ±$\pm 1$.
@@
The conormal cycles $[Y_{w}]$ span the Poincaré dual of middle-dimensional relative cohomology $H^{\dim_{\mathbb{C}}(X)}(X,X\setminus\mathbb{Y})$ of rank $|W|$.
@@
The conormal cycles $[Y_{w}]$ represent classes in the Borel–Moore homology of $\mathbb{Y}:=\bigcup_{w\in W}Y_{w}$, i.e. classes in $H_{*}^{\mathrm{BM}}(\mathbb{Y})\cong H^{*}(X,X\setminus\mathbb{Y})$ by Poincaré duality.
Choose a filtration of $\mathbb{Y}$ into locally closed subsets $\emptyset=Y_{0}\subseteq Y_{1}\subseteq\cdots\subseteq Y_{|W|}=\mathbb{Y}$ such that every $Y_{j}\setminus Y_{j-1}$ is a conormal of a Schubert cell $N^{*}\mathring{X}_{w}$. Note that the Borel–Moore homology $H_{*}^{\mathrm{BM}}(Y_{j}\setminus Y_{j-1})$ is of rank $1$ and concentrated in degree $\dim_{\mathbb{C}}(X)=\frac{1}{2}\dim_{\mathbb{R}}(X)$, spanned by the class of $[N^{*}\mathring{X}_{w}]$. Iterative application of the long exact sequence in Borel–Moore homology (see e.g. [CG97, (2.6.10)]) implies that $H_{*}^{\mathrm{BM}}(\mathbb{Y})\cong H^{*}(X,X\setminus\mathbb{Y})$ is concentrated in degree $\dim_{\mathbb{C}}(X)$ and is of rank $|W|$. Since for each $w\in W$, $Y_{w}=N^{*}\mathring{X}_{w}\cup\bigsqcup_{v<w}N^{*}\mathring{X}_{v}$, it follows that the Poincaré duals of $[Y_{w}]$ are linearly independent in $H^{*}(X,X\setminus\mathbb{Y})$, as desired. ∎
The equivariant cohomology classes in $H^{*}_{\mathbf{T}}(X)$ represented by the cycles $Y_{w}$ now factor through the map
(3.3) $H^{*}_{\mathbf{T}}(X,X\setminus\mathbb{Y})\to H^{*}_{\mathbf{T}}(X),$
in the long exact sequence of pairs.
@@
To define the stable bases, note that the cotangent fiber at any $T$-fixed point $wB/B\in G/B$ is a $T$-representation with ($T$-equivariant) Euler class $e_{w}=e^{T}(T^{*}_{wB/B}G/B)$. For $N_{w}=T_{(wB/B,0)}(X)$ the tangent space of $(wB/B,0)$ in $X=T^{*}(G/B)$, consider the $\mathbf{T}$-invariant decomposition $N_{w}=N_{w,+}\oplus N_{w,-}$ into $T$-weights that pair positively and negatively for our choice of dominant $\beta:S^{1}\to T$, respectively. (There are no 00-weights, as $\beta$ is chosen generically). By the duality from holomorphic symplectic form, one concludes that $N_{w,+}\cong N_{w,-}^{\vee}$ up to a twist of a character of the $S^{1}$-action rotating the holomorphic symplectic form.
Therefore, up to a sign, the $T$-equivariant Euler class of $N_{w}$ given by $e(N_{w})=e(N_{w,-})^{2}$ is a perfect square; from the polarization $T_{(wB/B,0)}(X)=T_{wB/B}G/B\oplus T^{*}_{wB/B}G/B$, we also have $e(N_{w})=e_{w}^{2}$. Now fix the sign ±$\pm e(N_{w,-})$ so that
@@
Below, we consider the isolated fixed point $(wB/B,0)\in X^{T}$ corresponding to a Weyl group element $w\in W$, and denote by $[w]\in H^{0}(X^{T})$ the corresponding fundamental class of the isolated point. For a class $\alpha\in H^{*}_{\mathbf{T}}(X)$, we define $\alpha|_{w}$ as the image of $\alpha$ under $H^{*}_{\mathbf{T}}(X)\to H^{*}_{\mathbf{T}}(X^{T})\cong\bigoplus H^{*}_{\mathbf{T}}(\mathrm{pt})\cdot[w]\to H^{*}_{\mathbf{T}}(\mathrm{pt})\cdot[w]$, i.e. restriction to the fixed locus followed by the projection to the summand generated (as a $H^{*}_{\mathbf{T}}(\mathrm{pt})$-module) by $[w]$.
@@
satisfying the following properties. (i) $\mathrm{Stab}_{+}(w)|_{H^{*}_{\mathbf{T}}(X\setminus\mathbb{Y})}=0$, (ii) $\mathrm{Stab}_{+}(w)|_{w}=\pm e(N_{w,-})$ where the sign is determined by (3.4), (iii) $\mathrm{Stab}_{+}(w)|_{v}$ is divisible by $h$ for $v<w$ in the Bruhat order.
@@
Property (i) with 3.7 implies that the stable envelope classes $\mathrm{Stab}_{+}(w)$ for $w\in W$ are linear combinations of the conormal cycles $Y_{w}$ in $H^{*}_{\mathbf{T}}(X)$. Property (ii) implies that the stable envelope classes form a basis in $H^{*}_{\mathbf{T}}(X)_{\mathrm{loc}}$ after localization; note that $e_{w}\neq 0$, as we are considering the Euler class of the normal bundle of the fixed locus (more precisely, its half).
Moreover, the stable envelopes are defined via Lagrangian correspondences ([MO19, Proposition 3.5.1]), and hence takes middle degree to middle degree. As $w\in W$ has a corresponding class $[w]:=[(wB/B,0)]\in H^{0}_{\mathbf{T}}(X^{T})$, the stable envelope class $\mathrm{Stab}_{+}(w)\in H^{\dim_{\mathbb{C}}X}_{\mathbf{T}}(X)$ lies in middle dimension (which is also evident from property (ii)). ∎
@@
Example 3.11.
In the case of $T^{*}\mathbb{P}^{1}$, the support of $\mathrm{Stab}_{+}$ are given by the (union of) Lagrangians given by (i) the cotangent fiber at the north pole for $1\in W$ and (ii) the union of the cotangent fiber at the north pole with the zero section $\mathbb{P}^{1}$ for $s\in W$.